# Broadway: Boost Your Text-to-Video Generation Model in a Training-free way

### v1

@@ 4.2 Fourier-based Motion Enhancement
where $\mathcal{F}$ denotes the FFT operation, $\mathbf{A}\in\mathbb{C}^{(B\times H\times W)\times F\times F}$ is the complex-valued matrix resulting from applying the FFT to $\mathcal{A}$, and $\tau$ is a hyperparameter that determines the frequency range for the high-pass and low-pass filters. As demonstrated in Fig. 6 (a), experiments involving the selective removal of high-frequency or low-frequency components from the temporal attention map during the denoising process have yielded insightful observations. Videos that retain only the low-frequency components tend to exhibit a nearly static structure, closely mirroring the characteristics of their unmodified counterparts. In contrast, videos that include solely high-frequency components display abundant motion but are marred by inconsistency and persistent flickering. These findings suggest that the essence of motion in generated videos is predominantly encapsulated within the high-frequency components of their temporal attention maps.
Figure 6: Frequency decomposition and manipulation. (a) Results obtained by directly removing either the high-frequency or low-frequency components from the temporal attention map. The motion in generated videos is primarily present in the high-frequency components of the temporal attention map. (b) Results obtained by scaling the high-frequency components by a factor of $\beta$.
@@
where $\widetilde{\mathcal{F}}$ represents the inverse Fast Fourier Transform (iFFT) operation, and $\mathcal{A}^{{}^{\prime}}$ signifies the temporal attention map with the scaled high-frequency components. Based on aforementioned equations, the following theorems can be proven. Please refer to Section A.1.1 for a comprehensive proof.
@@
As illustrated in Fig. 6 (b), setting $\beta=1.5$ amplifies the energy of the temporal attention maps, leading to greater motion magnitude. Conversely, setting $\beta=0.5$ results in reduced motion.
Figure 7: BroadWay Operations. (a) Temporal Self-Guidance. The temporal attention map from up_blocks.1 is injected into the corresponding modules of up_blocks.2/3 with a guidance ratio $\alpha$, in order to enhance the structural plausibility and temporal consistency. (b) Fourier-based Motion Enhancement. A scaling factor $\beta$ is applied to the high-frequency components of the temporal attention map, thereby amplifying the magnitude of the motion.
4.3 BroadWay
Leveraging Temporal Self-Guidance and Fourier-based Motion Enhancement, we introduce BroadWay, a parameter-free method that enhances the quality of text-to-video generation without increasing memory requirements or sampling time. As illustrated in Fig. 7, BroadWay initially applies Temporal Self-Guidance to improve the structural coherence and temporal consistency of the video. Subsequently, Fourier-based Motion Enhancement is employed to amplify motion dynamics. To ensure that the motion magnitude of generated videos processed by BroadWay exceeds that of the original, unenhanced videos, the energy of the temporal attention map after Fourier-based Motion Enhancement, denoted as $E_{3}$, must be greater than the energy of the original temporal attention map, denoted as $E_{1}$. To achieve this, the scaling factor $\beta$ is defined as a function of the energies before and after Temporal Self-Guidance, $E_{1}$ and $E_{2}$, respectively:
@@ 4.3 BroadWay
where $\beta_{0}$ is user-given value of $\beta$ to control the motion magnitude. $E_{2}^{H}$ and $E_{2}^{L}$ denoting the energies of the high-frequency and low-frequency of the attention map after applying Temporal Self-Guidance, respectively. Please refer to Section A.1.2 for a detailed proof for Eq. 7.
5 Experiments
Figure 8: Samples synthesized by AnimateDiff with or without BroadWay. The samples utilizing the BroadWay exhibit enhanced structural plausibility, temporal consistency, and an increased richness in motion dynamics.
@@ 5.1 Experiments Setup
Setting up. We mainly conduct our experiments on two mainstream diffusion based T2V backbones with superior visual quality: AnimateDiff ($512\times 512$(Guo et al., 2023b) and VideoCrafter2 ($320\times 512$(Chen et al., 2024). Results synthesized by vanilla T2V backbones are used as a benchmark. The BroadWay parameters are set to $\alpha=0.6$, $\beta=1.5$ in default for AnimateDiff, $\alpha=0.1$, $\beta=10$ in default for VideoCrafter2. BroadWay operations are only applied during the first 20% steps of the denoising process. DDIM sampler (Song et al., 2020) with classifier-free guidance (Ho & Salimans, 2022) is adopted in the inference phase.
Evaluation metrics. We report three metrics for quantitative evaluation. First, we conduct a user study with 30 participants to assess Video Quality, considering both structure coherence and motion magnitude. Secondly, we compare the Optical Flow values of 1000 videos generated by Vanilla T2V backbones and BroadWay-enhanced backbones. Additionally, we employ a multimodal large language model, GPT-4o (Achiam et al., 2023), for a comprehensive Multimodal-Large-Language-Model (MLLM) Assessment on hundreds of generated videos. Refer to Section A.3 for details.
@@ 5.2 Qualitative Comparison
As presented in Fig. 8 and Fig. 9, with the integration of BroadWay, various T2V backbones demonstrates a notable performance improvement compared to their vanilla synthesis results. For instance, giving AnimateDiff the prompt “a green wool doll is displayed on the wooden table.”, BroadWay enhances the structural consistency of the synthesized video, preventing the collapse of the doll’s head and tail. Moreover, in the “A jeep driving on the grass near a forest.” case, BroadWay amplifies the dynamic effects of the scene, making the jeep exhibit more pronounced motion. For VideoCrafter2, when provided with the prompt “A horse jumping over a fence during a race, crowd cheering.”, BroadWay reconstructs the structure of the rider and horse, addressing the issue of structural anomalies in the horse’s legs while enhancing the overall motion to appear more synchronized and aesthetically pleasing. In cases like “A penguin sliding on ice, snowy landscape in the background.”, BroadWay preserves the original structural integrity while introducing richer, more dynamic motion to the scene.
In summary, BroadWay effectively improves the structural consistency of synthesized videos while amplifying their motion dynamics, resulting in a significant enhancement in the overall synthesis quality of the T2V backbones.
Figure 9: Samples synthesized by VideoCrafter2 with or without BroadWay. The samples utilizing the BroadWay exhibit enhanced structural plausibility, temporal consistency, and an increased richness in motion dynamics.
@@ 5.3 Quantitative Evaluation
User Study. As shown in Table 1 (a), we present the voting results, expressed as percentages, for vanilla T2V backbones and BroadWay-enhanced backbones. Our analysis shows that BroadWay receives the majority of votes, demonstrating that BroadWay provides a substantial improvement to the T2V diffusion model in terms of overall video quality, taking into account both structure coherence and motion magnitude.
Motion Magnitude. To objectively evaluate the motion magnitude, RAFT (Teed & Deng, 2020) is introduced to estimate the forward optical flow between consecutive frames, and the average intensity value of estimated optical flow is used to quantify the motion magnitude within the video. As presented in Table 1 (b) BroadWay shows substantial improvements in mean motion intensity, indicates its efficacy in producing large-magnitude motion.
MLLM Assessment. In light of the impressive strides made by Multimodal-Large-Language-Models (MLLM) recently in image/video understanding, the state-of-the-art MLLM, i.e., GPT-4o (Achiam et al., 2023), is employed for video quality assessment, covering structural rationality and motion consistency. As can be observed in Table 1 (c)-(d), BroadWay exhibits notable gains in both metrics, validating its role in substantially improving overall video quality.
Table 1: Quantitative results with or without BroadWay.
Method (a) Video Quality (b) Optical Flow (c) Structural Rationality (d) Motion Consistency
AnimateDiff 25.42% 1.5743 41.94% 34.62%
+ BroadWay 74.58% 2.4673 58.06% 65.38%
VideoCrafter2 30.54% 1.5555 18.48% 39.60%
+ BroadWay 69.46% 3.6204 81.52% 60.40%
5.4 Ablation Study
Effects of Temporal Self-Guidance (TSG). Temporal Self-Guidance plays a critical role in reinforcing structural integrity, thus effectively mitigating structural breakdowns and preventing motion artifacts, as can be observed in Fig. 10 (a). However, it cannot enhance the magnitude of motion, showing limited improvement in scenarios with little motion, as shown in Fig. 10 (b).
Figure 10: Ablation study on BroadWay components. The left panel illustrates an instance of inconsistency artifacts present in the original video, whereas the right panel exhibits a scenario where the original video lacks sufficient motion.
Effects of Fourier-based Motion Enhancement (FME). Fourier-based Motion Enhancement is responsible for amplifying the motion dynamics in generated videos. In scenarios where the motion is insufficient, this technique effectively increases the dynamic content, as shown in Fig. 10 (b). However, motion enhancement alone does not guarantee appealing video quality when structural breakdown occurs, as illustrated in Fig. 10 (a).
Effects of BroadWay. By integrating Temporal Self-Guidance with Fourier-based Motion Enhancement, BroadWay achieves simultaneous enhancement of both the structural integrity and motion dynamics in generated videos (Fig. 10 (a)-(b) Ours vs. Vanilla).
@@ 5.5 Image-to-Video
Figure 11: Generated results by SparseCtrl with or without BroadWay.
Similar to text-to-video (T2V) tasks, image-to-video (I2V) is also a significant research area within video diffusion models. Here we employ SparseCtrl (Guo et al., 2023a), a strong and flexible structure control method, as the I2V backbone to preliminarily validate the potential of BroadWay in image-to-video tasks. As illustrated in Fig. 11, the infusion of BroadWay into SparseCtrl serves to enhance the dynamic effects of the synthesized video while preserving the structural integrity of the reference image. Specifically, we observe that the video synthesized with BroadWay exhibits more vivid wave motions, and the reflections of the setting sun display enhanced dynamic aesthetics. These experimental results demonstrate that BroadWay effectively enhances the quality of both T2V and I2V video generation tasks, positioning it as a versatile and powerful booster for video diffusion models.
@@ 6 Conclusion
In this study, we present BroadWay, a training-free method to improve the quality of text-to-video generation without introducing additional parameters, augmenting memory or sampling time. BroadWay is composed of Temporal Self-Guidance and Fourier-based Motion Enhancement. The former improves the structural plausibility and temporal consistency by reducing the disparity between the temporal attention maps across various decoder blocks. The later enhances the magnitude and richness of motion by scaling the high frequency of the temporal attention maps. The proposed method can be easily integrated with other T2V models in a plug-and-play manner, offering a general and effective solution to enhance video generation quality during inference phase.
Appendix A Appendix
@@
In the appendix, we present the proof of Fourier-based Motion Enhancement (Section A.1), additional qualitative results (Section A.2), details of our quantitative evaluation (Section A.3), as well as the limitations of our method(Section A.4), as a supplement to the main paper.
A.1 Proof of Fourier-based Motion Enhancement
@@
In this section, we provide a detailed proof of how Fourier-based Motion Enhancement alters the energy of the temporal attention map in the denoising process.
A.1.1 Frequency Manipulation

### v3

@@ 4.2.2 Motion Enhancement by Frequency Re-weighting
where $\mathcal{F}$ denotes the 1D FFT operation along the softmax axis, $\mathbf{A}\in\mathbb{C}^{(B\times H\times W)\times F\times F}$ is the complex-valued matrix resulting from applying the FFT to $\mathcal{A}$, and $\tau$ is a hyperparameter that determines the frequency range for the high-pass and low-pass filters. As demonstrated in Fig. 5, experiments involving the selective removal of high-frequency or low-frequency components from the temporal attention map during the denoising process have yielded insightful observations. Videos that retain only the low-frequency components tend to exhibit a nearly static structure, closely mirroring the characteristics of their unmodified counterparts. In contrast, videos that include solely high-frequency components display abundant motion but are marred by inconsistency and persistent flickering. These findings suggest that the essence of motion in generated videos is predominantly encapsulated within the high-frequency components of their temporal attention maps.
Figure 5: Frequency decomposition. By directly removing either the high-frequency or low-frequency components from the temporal attention map, it can be observed that motion in generated videos is primarily present in the high-frequency components.
@@
where $\widetilde{\mathcal{F}}$ represents the inverse Fast Fourier Transform (iFFT) operation, and $\mathcal{A}^{{}^{\prime}}$ signifies the temporal attention map with the scaled high-frequency components. Based on aforementioned equations, we have the following theorems (detailed proof is provided in the supplementary material).
@@
Subsequently, the motion magnitude of generated videos can be enhanced with improved energy under $\beta>1$.
Figure 6: ByTheWay Operations. (a) Temporal Self-Guidance. The temporal attention map from up_blocks.1 is injected into the corresponding modules of up_blocks.2/3 with a guidance ratio $\alpha$, in order to enhance the structural plausibility and temporal consistency. (b) Fourier-based Motion Enhancement. A scaling factor $\beta$ is applied to the high-frequency components of the temporal attention map, thereby amplifying the motion magnitude within generated videos.
4.3 ByTheWay Operations
Leveraging the techniques proposed above, we introduce ByTheWay, a training-free method to enhance the T2V quality without increasing inference expense. As illustrated in Fig. 6, ByTheWay initially applies Temporal Self-Guidance to improve the structural coherence and temporal consistency of the video. Subsequently, Fourier-based Motion Enhancement is employed to amplify motion dynamics. To ensure that the motion magnitude of generated videos processed by ByTheWay exceeds that of the original, unenhanced videos, the energy of the temporal attention map after Fourier-based Motion Enhancement, denoted as $E_{3}$, must be greater than the energy of the original temporal attention map, denoted as $E_{1}$. To achieve this, the scaling factor $\beta$ is defined as a function of the energies before and after Temporal Self-Guidance, $E_{1}$ and $E_{2}$, respectively:
@@ 4.3 ByTheWay Operations
where $\beta_{0}$ is user-given value of $\beta$ to control the motion magnitude. $E_{2}^{H}$ and $E_{2}^{L}$ denoting the energies of the high-frequency and low-frequency of the attention map after applying Temporal Self-Guidance, respectively. See the supplementary material for a detailed explanation for Eq. 7.
5 Experiments and Results
Figure 7: Samples generated by AnimateDiff[9] with or without ByTheWay.
Figure 8: Samples generated by VideoCrafter2[7] with or without ByTheWay.
@@ 5.1 Experiments Setup
Base models. We mainly conduct our experiments on two mainstream T2V backbones with superior visual quality: AnimateDiff ($512\times 512$[9] with Realistic Vision V5.1 LoRA and VideoCrafter2 ($320\times 512$[7]. Results generated by vanilla backbones are used as a baseline. ByTheWay operations are only applied during the first 20% steps of the denoising process. DDIM sampler [30] with classifier-free guidance [37] is adopted in the inference phase.
Parameter setup. The ByTheWay parameters are set to $\alpha=0.6$, $\beta=1.5$, $\tau=7$ in default for AnimateDiff, $\alpha=0.1$, $\beta=10$, $\tau=7$ in default for VideoCrafter2. Note that the default values of ByTheWay parameters are relatively robust within a specific T2V backbone but may not be universally optimal for different backbones, since different base models exhibit variations in their motion preferences.
Evaluation metrics. We report three metrics for quantitative evaluation. First, we conduct a user study with 30 participants to assess Video Quality, considering both structure coherence and motion magnitude. Secondly, we employ GPT-4o [60] for a comprehensive Multimodal-Large-Language-Model (MLLM) Assessment on hundreds of generated videos. The implementation details are available in the supplementary material. Moreover, we evaluate 200 videos generated by Vanilla T2V backbones and ByTheWay-enhanced backbones using VBench [70].
@@ 5.2 Qualitative Comparison
As presented in Fig. 7 and Fig. 8, with the integration of ByTheWay, various T2V backbones demonstrates a notable performance improvement compared to their vanilla synthesis results. For instance, giving AnimateDiff the prompt “a green wool doll is displayed on the wooden table.”, ByTheWay enhances the structural consistency of the synthesized video, preventing the collapse of the doll’s head and tail. Moreover, in the “A jeep driving on the grass near a forest.” case, ByTheWay amplifies the dynamic effects of the scene, making the jeep exhibit more pronounced motion. For VideoCrafter2, when provided with the prompt “A horse jumping over a fence during a race, crowd cheering.”, ByTheWay reconstructs the structure of the rider and horse, addressing the issue of structural anomalies in the horse’s legs while enhancing the overall motion to appear more synchronized and aesthetically pleasing. In cases like “A penguin sliding on ice, snowy landscape in the background.”, ByTheWay preserves the original structural integrity while introducing richer, more dynamic motion to the scene.
FreeInit [71] is a training-free method designed to improve T2V temporal consistency by iteratively refining the spatial-temporal low-frequency components of the initial latent throughout the denoising process. As shown in Fig. 9, regardless of whether the vanilla-generated video is corrupted, FreeInit leads to a significant loss of motion while refining its structure, resulting in a nearly static video. In contrast, ByTheWay simultaneously enhances both structural coherence and motion magnitude, achieving a comprehensive improvement in video quality. Additionally, FreeInit requires iterative cycles to refine the initial noise, adding significant time overhead (5$\times$ sampling time), while ByTheWay incurs almost no extra inference cost.
In summary, ByTheWay effectively improves the structural consistency of synthesized videos while amplifying their motion dynamics, resulting in a significant enhancement in the overall synthesis quality of the T2V backbones.
Figure 9: Visual comparison with FreeInit [71].
@@ 5.3 Quantitative Evaluation
User Study. As shown in Table 1 (a), ByTheWay receives the majority of votes, demonstrating its remarkable improvement in generating visually appealing results.
Evaluation Metric (a) User Study (b) MLLM Assessment
Video Quality Structure Rationality Motion Consistency
AnimateDiff [9] 25.42% 41.94% 34.62%
+ ByTheWay 74.58% 58.06% 65.38%
VideoCrafter2 [7] 30.54% 18.48% 39.60%
+ ByTheWay 69.46% 81.52% 60.40%
Table 1: Voting results of user study and MLLM assessment.
Evaluation Metric VBench [70] Metrics
Subject Consistency $\uparrow$ Background Consistency $\uparrow$ Motion Smoothness $\uparrow$ Dynamic Degree $\uparrow$ Aesthetic Quality $\uparrow$ Imaging Quality $\uparrow$
AnimateDiff [9] 0.9318 0.9507 0.9474 0.4073 0.6376 0.7215
AnimateDiff + FreeInit [71] 0.9712 0.9695 0.9713 0.2941 0.6571 0.7459
AnimateDiff + ByTheWay 0.9744 0.9725 0.9786 0.5245 0.6609 0.7553
VideoCrafter2 [7] 0.9732 0.9691 0.9749 0.4086 0.6477 0.6575
VideoCrafter2 + ByTheWay 0.9852 0.9896 0.9863 0.5547 0.6598 0.6738
Table 2: Quantitative results of ByTheWay on VBench [70]. ByTheWay facilitates the best performance of different T2V models.
Figure 10: Ablation on ByTheWay parameters. Left: “close up photo of a rabbit, forest, haze, …”; Middle: “car runs in the forest, …”; Right: “Kid with curly hair plays in the park, …”. Dashed boxes indicates the optimal parameters we chose in the experiments.
MLLM Assessment. In light of the impressive strides made by Multimodal-Large-Language-Models (MLLM) recently in image/video understanding, the state-of-the-art MLLM, i.e., GPT-4o [60], is employed for video quality assessment, covering structural rationality and motion consistency. As shown in Table 1(b), ByTheWay exhibits notable gains in both structure rationality and motion consistency, validating its role in substantial video quality enhancement.
VBench Metrics. To objectively evaluate the overall generation quality, VBench [70] is introduced to serve as a comprehensive benchmark. As presented in Table 2, ByTheWay shows substantial improvements in VBench metrics, indicates its efficacy in enhancing T2V generation quality. Moreover, it can be observed that ByTheWay outperforms FreeInit [71], which is a strong training-free solution, across all evaluated dimensions.
@@ 5.4 Image-to-Video
Similar to text-to-video (T2V) tasks, image-to-video (I2V) is also a significant research area within video diffusion models. Here we employ SparseCtrl [42], a strong and flexible structure control method, as the I2V backbone to preliminarily validate the potential of ByTheWay in image-to-video tasks. As illustrated in Fig. 11, the infusion of ByTheWay into SparseCtrl serves to enhance the dynamic effects of the synthesized video while preserving the structural integrity of the reference image. Specifically, we observe that the video synthesized with ByTheWay exhibits more vivid wave motions, and the reflections of the setting sun display enhanced dynamic aesthetics.
These experimental results demonstrate that ByTheWay effectively enhances the quality of both T2V and I2V generation tasks, positioning it as a versatile and powerful booster for video diffusion models.
Figure 11: Results of SparseCtrl [42] with/without ByTheWay.
5.5 Ablation Study and Analysis
Effect of $\alpha$. $\alpha$ represents the infusion ratio of lower-level temporal attention information in Temporal Self-Guidance. As shown in Fig. 10, an appropriate $\alpha$ strengthens the temporal consistency of the video, but an excessively large $\alpha$ may lead to the loss of motion information.
Effect of $\beta$. $\beta$ stands for the scaling factor of the high-frequency components in the temporal attention map within Fourier-based Motion Enhancement. An appropriate $\beta$ introduces richer and more intensified motion to the video, but an excessively large $\beta$ may cause the emergence of unexpected motion artifacts, as can be observed in Fig. 10.
Effect of $\tau$. $\tau$ denotes the number of discrete frequency components involved in Fourier-based Motion Enhancement. As shown in Fig. 10, a larger $\tau$ allows for the manipulation of more frequency components that encode motion, thus promoting the motion enhancement effect.
@@ 6 Conclusion
In this work, we present ByTheWay, a training-free method to improve the quality of video generation without introducing additional parameters, augmenting memory or sampling time. ByTheWay consists of two key components: Temporal Self-Guidance and Fourier-based Motion Enhancement. The former improves structural plausibility and temporal consistency by reducing the disparity between temporal attention maps of different hierarchical levels. The latter enhances motion magnitude by scaling the high frequency of temporal attention maps. The proposed method can be easily integrated with available T2V backbones in a plug-and-play manner, offering a general and effective solution to enhance video generation quality during the inference phase.
Supplementary Material
@@
In the supplementary material, we present additional qualitative results (Section 7), more ablation experiments (Section 8), details of our user study and MLLM assessment (Section 9), the proof of Fourier-based Motion Enhancement (Section 10), as well as the limitation of our method (Section 11), as a supplement to the main paper.
7 Additional Qualitative Results
More Results on AnimateDiff. We present more results for video motion enhancement (Fig. 17) and structure enhancement (Fig. 18) on AnimateDiff.
More Results on VideoCrafter2. We present more results for video motion enhancement (Fig. 19) and structure enhancement (Fig. 20) on VideoCrafter2.
Application on DiT-based Architecture. We extend ByTheWay to a DiT-based T2V backbone, CogVideoX. As depicted in Fig. 13, ByTheWay showcases potential for motion enhancement in diffusion DiT.
8 Additional Ablation Study
Choice of Guidance Anchor. Fig. 14 demonstrates that up_blocks.1 is the bottleneck in video motion modeling, injecting the temporal attention map of up_blocks.1 into subsequent decoder blocks helps align motion modeling across different levels of diffusion U-Net, thus enhance temporal consistency. In contrast, injecting information from later blocks fails to achieve this goal. Note that using up_blocks.3 as the anchor implies the absence of Temporal Self-Guidance (vanilla result).
Number of Operation Steps. Fig. 16 reveals that the initial $20\%$ of sampling steps play a crucial role in shaping video motion, making the application of ByTheWay operations beyond this point have minimal effect on the generation quality. Moreover, when ByTheWay operations are applied only during $20\%$ to $80\%$ of the sampling steps, the generated video appears almost identical to the original video, which can be attributed that video motion is mainly determined by the early denoising stage.
Does More Sampling Steps Help? As shown in Fig. 15, the vanilla T2V backbone with 5×5$\times$5 × sampling steps is inferior to incorporating the ByTheWay-enhanced backbone with only 1×1$\times$1 × sampling steps, this demonstrates that ByTheWay is not equivalent to simply increasing the DDIM sampling steps.
9 Details of User Study & MLLM Assessment
User Study Details. In our user study, each participant receives 50 videos synthesized by Vanilla T2V backbones and 50 videos synthesized by ByTheWay-enhanced backbones. These videos are sampled from the same random seeds to ensure fair comparison. For each video pair from Vanilla and Vanilla + ByTheWay, participants are required to select the video they perceive as superior based on overall Video Quality, considering both structure coherence and motion magnitude, and cast their vote accordingly. The videos were presented in a randomized order to reduce potential bias, and participants were allowed ample time to review each pair before making their selections.
MLLM Assessment Prompt. Here, we present the prompt used in the MLLM assessment.
You are provided with two sets of video frames, each containing 4 representative frames, along with a shared textual prompt that was used to generate both videos. Your task is to perform a comparative evaluation of the two videos, focusing on their structure rationality / motion consistency.
Here is the frame data of Video_1.
Here is the frame data of Video_2.
Based on your evaluation of motion consistency, choose the video set you find to be superior. If you determine that the first set of frames (Video_1) is better, respond with ”A”. If the second set (Video_2) is superior, respond with ”B”. Return only ”A” or ”B” based on your assessment.
10 Fourier-based Motion Enhancement Proof
@@
In this section, we provide a detailed proof of how Fourier-based Motion Enhancement alters the energy of the temporal attention map in ByTheWay operations.
10.1 Frequency Components Manipulation